What is the amount of work done when two moles of ideal gas is compressed from a volume of I m' to 10 dm at 300 K against a pressure of 100 kPa
1
Expert's answer
2018-11-28T13:56:47-0500

m^3 = 1000 dm^3. Everyone knows that! 10 dm^3 = 0.01 m^3. W = -PdV = -P * (V2 - V1) = -10^5 (Pa) * (1 (m^3) - 0.01 (m^3)) = -99 (kJ).